Gil. Is reading a book that has 276 pages he has already read some of it last week he plans to read 20 pages tomorrow by then he will be 2/3 of the way through the book how many pages did gil read last week?
Answer by richwmiller(17219)

x+20=2/3*276
x+20 = 184
x=164